EXERCISING ESS 3. Calculate column percentages for Table W8.1: Table W8.1u Welfare Outcomes by Level of Education Education Welfare outcome No H.S. degree Low Medium High Note: HS. = high school H.S. degree 60 35 25 Some college 10 15 20 30 Note: HS. = high school 4. Practice writing up the results for the preceding question Identify the number of recipients and how they differ by education and welfare outcome. Then discuss the relationships between these two variables. Use statements in the form of Whereas xx percent of recipients with no high school degree have high welfare outcomes, xx percent of recipients with some college have high welfare outcomes. and so on.

#4
Answer #1

#3 -

Education
Welfare outcome No. H.S. degree H.S degree Some college Row Total Row %
Low 60 55 10 125 41%
Medium 35 55 15 105 34%
High 25 30 20 75 25%
Col. Total 120 140 45 305
Col. % 39% 46% 15%

Step 1- Calculate column total for each column .Example for first col.(No. H.S. degree) is (60+35+25) = 120

Step 2 - Aggregate column total as (120+140+45) = 305

Step 3 - Divide column total by aggregated column total and multiply with 100 .Example for 'No. H.S. degree , 120/305 * 100 =39%

# 4 -

There are total 305 recipients.

To discuss relationship between two variables, we'll take help of the below table where we are calculating percentage of each cell with respect to column total.As for example, the first cells says 50% of no high school degree holders having low welfare.

Education
Welfare outcome No. H.S. degree H.S degree Some college
Low 50% 39% 22%
Medium 29% 39% 33%
High 21% 21% 44%

- Out of all people with 'No. H.S. degree', most of the people (around 50%) are having 'Low' welfare outcome and only 21% of them are having 'High' welfare outcome.

- People with 'H.S. degree' are having almost same(39%) 'Low' and 'Medium' welfare outcome and low(21%) 'High' welfare outcome

- where as people having 'Some college' degree, are having 'High' welfare outcome around 44% and 'Low' welfare outcome around 22%

Thus from the above three observations, we can say the Welfare outcome is closely associated with the education level of a person Higher the education more likely the person would have higher welfare outcome.